how do i solve the following, 5+|3n-1|=2 ?
Answer by ptth176(50)

5+|3n-1|=2
=>|3n-1|=2-5
=>|3n-1|=-3
=>3n-1=-3 or 3n-1=-(-3)
=>3n=-2 or 3n=4
=>n=-2/3 or n=4/3